Your annual salary at your job is $49,843, and must pay 17% of your income in taxes. You then decide to save 9% of your disposable income. How much do you have left for consumption spending each month?

Round to two decimal places.
Answer #1

Solution:

Annual salary = $49,843

Disposable income = income - taxes

Disposable income = 49843 - 17%*49843

Disposable income = 49843(1 - 0.17) = $41,369.69

Consumption = disposable income - savings

Consumption = disposable income - 9%*disposable income or disposable income (1 - 9%)

Consumption = 41,369.69*(1 - 0.09) = $37,646.4179

This is the amount left for consumption spending annually, so for each month = 37,646.4179/12 = 3137.20149

So, amount left for consumption spending each month is $3,137.20 approximately.